OpenStudy (bekkah323):

Rewrite without rational exponents, and simplify, if possible. 49^3/2

OpenStudy (anonymous):

with a rational exponent, the numerator is the power and the denominator is the root

OpenStudy (anonymous):

making this \[\sqrt{49^3}\] take the square root first, cube second

OpenStudy (anonymous):

so i guess it would be easier if you write \(\sqrt{49}^3\)

OpenStudy (bekkah323):

where did the 2 go? I'm sorry i suck at radical problems!

OpenStudy (anonymous):

the two in the denominator means take the "square root"

OpenStudy (anonymous):

notice that the problem says exactly "Rewrite without rational exponents" when you get rid of the rational exponent, the 2 in the denominator becomes the square root sign

OpenStudy (bekkah323):

so since 7 times 7 is 49 would i then do 49^3?

OpenStudy (anonymous):

you could write \[49^{\frac{3}{2}}=\sqrt[2]{49^3}\] but no one writes an "index" of two

OpenStudy (bekkah323):

or 7^3

OpenStudy (bekkah323):

ahhh ok

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

as you said, \(7\times 7=49\) which means \(\sqrt{49}=7\)

OpenStudy (bekkah323):

so 343 is the answer?

OpenStudy (anonymous):

therefore \[\sqrt{49}^3=7^3=343\] yes you are right

OpenStudy (bekkah323):

ahhh ok, that makes sense! thank you!

OpenStudy (anonymous):

yw you got more or is that it?

OpenStudy (bekkah323):

I have a lot more problems with radicals!!!! One of my other questions says Rewrite with positive exponents, and simplify if possible? would i do the same thing one of the problems is 27^-1/3

OpenStudy (anonymous):

this one is a little different because of the minus sign in the exponent

OpenStudy (anonymous):

\[\large 27^{\frac{1}{3}}=\sqrt[3]{27}\] see this time you write the index because of the cubed root do you know what the cubed root of 27 is?

OpenStudy (bekkah323):

3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

now the minus sign in the exponent does not mean make it negative, it means take the reciprocal, in other words, flip it do you know what the reciprocal of 3 is ( don't think too hard)

OpenStudy (bekkah323):

-3?

OpenStudy (bekkah323):

or 1/3

OpenStudy (anonymous):

nah you fell for it don't make it negative take the reciprocal

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

so \[\large 27^{-\frac{1}{3}}=\frac{1}{\sqrt[3]{27}}=\frac{1}{3}\]is the right answer

OpenStudy (anonymous):

still need help?

OpenStudy (bekkah323):

that one makes sense, can you help me with a different question?

OpenStudy (anonymous):

yes of course

OpenStudy (bekkah323):

thanks, another questions says use the laws of exponents to simplify. Write the answers with positive exponents. 5^3/4 * 51/8

OpenStudy (anonymous):

i take it this is \[\large 5^{\frac{3}{4}}\times 5^{\frac{1}{8}}\] right?

OpenStudy (anonymous):

your job here is only to add \(\frac{3}{4}+\frac{1}{8}\) to get the exponent

OpenStudy (bekkah323):

yes, and 7/8

OpenStudy (anonymous):

you are right since \[\frac{3}{4}+\frac{1}{8}=\frac{6}{8}+\frac{1}{8}=\frac{7}{8}\] your answer is \[\large 5^{\frac{7}{8}}\]

OpenStudy (bekkah323):

thats it? thats easy!

OpenStudy (anonymous):

yeah adding exponents is pretty easy if you can add some people get stuck at adding fractions any more?

OpenStudy (bekkah323):

haha yah! if thats ok?

OpenStudy (anonymous):

yeah sure

OpenStudy (bekkah323):

it says use rational exponents to simplify. write the answer in radical notation if appropriate. \[(\sqrt[3]{ab})\]^15

OpenStudy (bekkah323):

sorry the ^15 is supposed to be right next to the equation

OpenStudy (anonymous):

\[\large\left(\sqrt[3]{ab}\right)^{15}\] right?

OpenStudy (bekkah323):

yup

OpenStudy (anonymous):

you are going to be amazed (maybe) at how easy this one is how many times does 3 go in to 15?

OpenStudy (bekkah323):

5 times

OpenStudy (anonymous):

ok so the answer is \(\left(ab\right)^5\)

OpenStudy (bekkah323):

wow that is easy! thanks

OpenStudy (anonymous):

don't forget you can go backwards too \[\sqrt[3]{x}=x^{\frac{1}{3}}\] so \[\large \sqrt[3]{x}^{15}=x^{\frac{15}{3}}=x^5\]

OpenStudy (anonymous):

yeah and at the beginning you said you sucked at radicals if i recall...

OpenStudy (anonymous):

any more of these?

OpenStudy (bekkah323):

haha yah, i just get confused by the different questions i guess!

OpenStudy (bekkah323):

nd yah, but i have to go back and apply this to some other questions first, so im good for a little bit! thanks!!

OpenStudy (anonymous):

yw good luck, hope this helped